OpenStudy (anonymous):

an equation that defines y as a function f of x is given a.) solve the equation for y in terms of x, and replace y with with the function f(x) b.)find f(4) 2x+7y=23

OpenStudy (anonymous):

2x+7y = 23 -> y = - 2x/7 + 23/7 => f ( 4) = -2*4/7 + 23/7 = 15/7